Imagine that Linda is studying health habits. She randomly surveys her friends on social media and asks each person how many times he or she exercises in one month. A total of 50 friends respond to Lindas survey, reporting the number of times they exercised in one month as a whole number. Linda displays the data she has collected in this frequency histogram 14 12 10 0 0 2345 678910 11 Sessions per month Find the median number of times exercised in one month as determined by Lindas survey times per month median
Answer #1
sessions frequency cumulative freq
0 7 7
1 10 17
2 9 26
3 4 30
4 1 31
5 2 33
6 0 33
7 0 33
8 11 44
9 4 48
10 2 50
50

from above middle value =(n+1)/2 th value =(50+1)/2 th value =25.5 th value =2

therefore median =2 times per month